OpenStudy (anonymous):

Negative 2 times sqr. root of 7 = ? Sqr. root of negative 14 or negative sqr. root of 14?

ganeshie8 (ganeshie8):

\(\large -2 \times \sqrt{7}\)

OpenStudy (anonymous):

You can't simplify it?

ganeshie8 (ganeshie8):

its already in simplified form, but if u want to put everything under radical, here is how we do it :- \(\large -1 \times 2 \times \sqrt{7}\)

ganeshie8 (ganeshie8):

\(\large -1 \times \sqrt{2} \times \sqrt{2} \times \sqrt{7}\)

ganeshie8 (ganeshie8):

\(\large -1 \times \sqrt{2 \times 2 \times 7}\)

ganeshie8 (ganeshie8):

\(\large -\sqrt{28 }\)

ganeshie8 (ganeshie8):

leave the - outside

OpenStudy (anonymous):

Ohh, thanks!!

ganeshie8 (ganeshie8):

np :) we leave - outside cuz, when u square you always get a positive real number. we never get a negative real number when we square.